OpenStudy (anonymous):

Solve the system by elimination: 5x + 4y = 12 3x - 3y = 18. Help? I'm stuck :/

OpenStudy (lucaz):

3x-3y=18 same equation: x-y=6 multiply by 4: 4x-4y=24 then subtract from the 1st equation and solve for x